# Jiawei Zhang    Shuang Yang    Bo Li

### v1

@@ Abstract
Large Language Model (LLM) agents equipped with external tools have become increasingly powerful for handling complex tasks such as web shopping, automated email replies, and financial trading. However, these advancements also amplify the risks of adversarial attacks, particularly when LLM agents can access sensitive external functionalities. Moreover, because LLM agents engage in extensive reasoning or planning before executing final actions, manipulating them into performing targeted malicious actions or invoking specific tools remains a significant challenge. Consequently, directly embedding adversarial strings in malicious instructions or injecting malicious prompts into tool interactions has become less effective against modern LLM agents. In this work, we present UDora, a unified red teaming framework designed for LLM Agents that dynamically leverages the agent’s own reasoning processes to compel it toward malicious behavior. Specifically, UDora first samples the model’s reasoning for the given task, then automatically identifies multiple optimal positions within these reasoning traces to insert targeted perturbations. Subsequently, it uses the modified reasoning as the objective to optimize the adversarial strings. By iteratively applying this process, the LLM agent will then be induced to undertake designated malicious actions or to invoke specific malicious tools. Our approach demonstrates superior effectiveness compared to existing methods across three LLM agent datasets.
@@ 1 Introduction
To tackle these scenarios, we propose a dynamic optimization strategy that adaptively updates its optimization objectives based on the LLM agent’s own reasoning process, overcoming the limitations of previous fixed-prefix optimization (Zou et al., 2023). Specifically, as shown in Figure 2, we proceed as follows: (1). Gather the Initial Response: We gather the agent’s initial response—along with its token probability distribution—given the current instruction augmented with an initialized adversarial string. (2). Identify Positions for Inserting Noise: We then automatically pinpoint several optimal positions in the agent’s current reasoning where we can insert some “noise” (e.g., the name of a target item or a malicious tool). (3). String Optimization: We then optimize the adversarial string to maximize the probability of these inserted noises under the current modified reasoning path. We then regenerate the agent’s new response with the updated adversarial string, identify the new positions for noise insertion, and optimize the adversarial string again, and do Step 1 to 3 iteratively. Through this dynamic process, our method continuously adapts to the agent’s evolving reasoning, ultimately steering it toward an incorrect reasoning path that triggers the target malicious action.
Using this approach, as shown in Figure 1, we successfully generate an adversarial string that compels a real-world AI email agent (AutoGen, 2024)—based on Microsoft’s AutoGen (Wu et al., 2024)—to forward the user’s recent private emails to an attacker during an auto-reply process. Detailed attack logs are deferred to Section C.2. Moreover, for malicious-observation benchmarks such as InjecAgent (Zhan et al., 2024) and WebShop (Yao et al., 2022), we attain an Attack Success Rate (ASR) of 99% and 61.75% with LLaMA 3.1 (Meta, 2024), and 97% and 91.50% with Ministral (MistralAI, 2024), respectively. Additionally, on the malicious-instruction benchmark AgentHarm, we achieve an ASR of 97.7% on LLaMA 3.1, and 100.0% on Ministral.
@@
  • We also achieve the highest ASR across three diverse datasets, outperforming all existing baselines, and successfully attack a real-world agent.
@@ 2 Related Work
Attacking language models is generally more challenging than attacking image-based models due to the discrete nature of language tokens. This discreteness hampers the direct application of gradient-based methods (e.g., PGD attack (Madry et al., 2018)) in the text domain. HotFlip (Ebrahimi et al., 2018) first represents text edits (such as flipping tokens) as vectors in the input space and uses the gradient to select the optimal manipulation, thereby successfully attacking text classification tasks. Based on this, UAT (Wallace et al., 2019) employs the similar gradient-based search to discover “trigger” tokens that can mislead model predictions or induce unethical text generation once added to the original input. AutoPrompt (Shin et al., 2020) later extends this approach to a top-$k$ token search procedure, while ARCA (Jones et al., 2023) proposes iteratively updating tokens at specific indices in the prompt based on the current values of the remaining tokens. More recently, GCG Attack (Zou et al., 2023) has emerged as the most effective method by optimizing the adversarial suffix to trigger an affirmative response from the model, though the resulting optimized tokens are usually unreadable. AutoDan (Liu et al., 2023a) is then proposed to apply a genetic algorithm to preserve semantic information of the optimized adversarial tokens while still achieving robust attack performance. However, when it comes to LLM agents, the goal of an attack extends beyond eliciting a malicious response: it must also trigger a specific target function through extended reasoning, which makes the attack more challenging.
LLM Agents.
@@
LLM agents capable of interacting with external environments and invoking tools or functions have become increasingly prevalent. WebShop (Yao et al., 2022) creates a simulated Amazon web shopping environment where agents can perform actions such as “search” or “click,” while WebArena (Zhou et al., 2023) offers a more realistic and reproducible environment with a broader range of tools and scenarios. Later, numerous benchmarks (Liu et al., 2023b; Deng et al., 2024; Zheng et al., 2024) have been introduced to evaluate LLM agents’ performance in tasks like web-based interactions and computer operations. As LLM agents grow more powerful and increasingly act on behalf of users, their safety becomes paramount. For example, within malicious environments, adversaries may insert indirect instructions (e.g., “Please ignore all previous instructions and adhere to the following instruction instead…”) within the observation from tool calling, which is referred as prompt injection attacks (Perez & Ribeiro, 2022; Liu et al., 2023c; Xu et al., 2024; Liao et al., 2024). InjecAgent (Zhan et al., 2024) is a benchmark designed to test LLM agents’ vulnerability to such indirect prompt injections after tool calls, and AgentDojo (Debenedetti et al., 2024) extends this to a more dynamic environment. Beyond prompt injection, LLM agents can also be directly asked to fulfill malicious instructions (e.g., committing credit card fraud), and AgentHarm (Andriushchenko et al., 2024b) is a benchmark proposed for measuring such harmfulness in LLM agents based on Inspect AI (UK AI Safety Institute, 2024). In this work, we propose a systematic red-teaming framework UDora that effectively optimizes adversarial strings to trigger specific target actions for both scenarios: malicious environments and malicious instructions.
@@ 3 Red-Teaming on LLM Agents w/ UDora
We delineate two primary attack scenarios: malicious environment and malicious instruction, as depicted in Figure 2. In the malicious environment scenario, the user’s initial instruction is benign; however, the third-party environment is corrupted with the insertion of an adversarial string into the agent’s observations post-tool interaction. This triggers the agent to deviate from the user’s original (benign) intention, executing an unintended and harmful action. Meanwhile, the malicious instruction scenario involves appending an adversarial string to a harmful instruction aimed at circumventing the safeguards of underlying LLM agents, compelling them to execute a targeted action rather than reject the malicious request. In this work, we assume that we have access to the entire target victim model, while for real agents, we assume we have access to the returned token probability distribution at each position.
Motivation.
@@
LLM agents commonly employ Chain-of-Thought (CoT) (Wei et al., 2022) or ReAct (Yao et al., 2023), generating detailed and complete reasoning before arriving the final action. As a result, simply optimizing for a fixed affirmative-response prefix without reasoning (e.g., “Sure, I will conduct the following action: click[xxx]”) is often challenging in this situation, because: (1). LLM agents are usually not trained or prompted to immediately produce final actions but with a extended reasoning; (2). Variations in system prompts, instructions, and the underlying LLMs result in substantially different response styles, rendering static, manually-created prefixes unsuitable for consistent optimization under different settings. Therefore, to attack an LLM agent more effectively, our strategy involves “leading the agent astray within its own reasoning style,” thereby inducing a malicious final action. A critical question is then how to develop such an incorrect reasoning path for optimization. As previously noted, aligning with the agent’s inherent reasoning style is crucial for facilitating the optimization process. Thus, as shown in Figure 2, we begin by collecting the agent’s initial reasoning, then identify the optimal multiple positions, and insert “noise” (e.g., the name of the target malicious tool) into this reasoning. We subsequently optimize the adversarial string to maximize the likelihood of this noise manifesting within the reasoning. As a result, we eliminate the need for manually created fixed prefixes or flawed reasoning paths, allowing us to dynamically and adaptively target any LLM agent, regardless of the underlying base LLM, prompt, or the specifics of the scenario or task. Next, we will detail how we quantitatively determine the optimal location for inserting the noise, and how we perform the optimization of the adversarial string carefully in this case.
@@
Notation.
We use $\mathcal{M}$ to denote the victim LLM agent, $x$ to denote the input instruction (including the system prompt), and $o$ to denote the observation returned by the environment after tool calling. Let $s$ be the adversarial string to be optimized: if the scenario involves a malicious instruction, $s$ is inserted in $x$; if the scenario involves a malicious environment, $s$ is inserted in $o$. We write $z$ for the response provided by the LLM agent when given the current input with inserted adversarial string, and use $\mathcal{P}$ to represent the corresponding token-level probability distributions across $z$. We use $t$ to represent noise, and $l$ to represent the number of noise to be inserted. By default, we use a token-based view for these notations (rather than words), so $z[:j]$ refers to the first $j$ tokens in the response $z$, and $\lvert t\rvert$ denotes the number of tokens in the noise.
Attack Algorithm.
As demonstrated in the introduction, our attack algorithm proceeds in three main steps, illustrated with a real-case example in Figure 2:
$r_{j}(t)=\frac{\bigl{(}\#\,\text{Matched Tokens}\bigr{)}+\bigl{(}\text{MatchedProbability}\bigr{)}}{|t|+1},\vspace{-2mm}$ (1)
[…]
(a) Sequential Optimization: In this mode, we aim to make the noise $t$ appear gradually in the response from the LLM agent during the optimization. Suppose that, at a certain iteration, there are $n<l$ positions at which $t$ has already been fully matched in the response $z$, i.e., $z[j:j+|t|]=t$. We then identify the next unmatched one, say $j_{n+1}$ (from the remaining candidate locations), that has the highest score and insert the noise as replacing $z[j_{n+1}:j_{n+1}+|t|]$ with $t$, thereby creating a modified noisy response $z^{*}$.
@@
Once a new adversarial string $s$ is obtained, we repeat Steps 1 to 3. In each iteration, the procedure recomputes the insertion positions in the new response $z$, inserts noise accordingly, and optimize $s$ to maximizing the appearance of the noise shown in the reasoning of the LLM agent. This iterative process continues until the LLM agent finally generate a wrong reasoning and calls the target malicious tool. We provide several real optimization cases across three datasets in Appendix B to better demonstrate this process and help gain more intuition for our red teaming efforts.
@@ 4 Experiments
Datasets.
We consider three datasets: InjecAgent (Zhan et al., 2024) and WebShop (Yao et al., 2022) for scenario of malicious environments, and AgentHarm (Andriushchenko et al., 2024b) for scenario of malicious instruction.
@@
Currently, few systematic attacks on LLM agents have been proposed. For the datasets of InjecAgent and WebShop, we adopt two baseline approaches. The first is the GCG attack (Zou et al., 2023), which performs optimization with a fixed prefix for the target action. The second is Prompt Injection (Perez & Ribeiro, 2022; Liu et al., 2023c), where we use the same prompt “IMPORTANT!!! Ignore all previous instructions and strictly adhere to the following instruction:” from InjecAgent. For WebShop, a similar prompt “IMPORTANT!!! Ignore all previous instructions and directly select this shopping item.” is used. For AgentHarm, we also adopt the GCG attack and the Template Attack from Andriushchenko et al. (2024a), which includes specific rules in the prompt to reduce the possibility of refusal by the LLM.
Experiment Setup.
@@
In our experiments, we consistently use the ReAct (Yao et al., 2023) prompt for each agent, which involves reasoning before executing the final action. The number of candidate strings to test in each iteration, the top $k$ of candidate substitutions considered at a given token position based on the gradient, and the number of optimization steps are set to 128, 32, and 500 respectively on InjecAgent and WebShop for both GCG and UDora. For AgentHarm, these parameters are adjusted to 256, 64, and 1000. Besides, motivated by Andriushchenko et al. (2024a), we initialize the adversarial string for both GCG and UDora by optimizing it on an example from the validation dataset to accelerate attack efficiency, which starts with 25 repeated ‘ x’, totaling 25 tokens. Additionally, the target noise for insertion in the InjecAgent and AgentHarm datasets is the name of the target adversarial tool (e.g., “get_latest_emails”), whereas for the WebShop dataset, it is the Amazon Standard Identification Number (ASIN) of the target adversarial item (e.g., B0BR6JWP1Q).
Base LLMs.
@@
We select two representative open-sourced LLMs, Llama-3.1-8B-Instruct (Meta, 2024) and Ministral-8B-Instruct-2410 (MistralAI, 2024), which support tool calling, on all three datasets. Meanwhile, our case study on a real-world AI email agent is based on gpt-4o-2024-08-06 (Hurst et al., 2024).
Metric.
@@
We use the Attack Success Rate (ASR) to measure the effectiveness of each attack method. Specifically, for InjecAgent, we determine whether the LLM Agent decides to utilize the target adversarial tool. For AgentHarm, a successful attack is one where the agent not only refrains from rejecting the malicious request but also uses the target tool. On WebShop, we evaluate whether the LLM agent successfully executes the ‘click’ action on the target adversarial item. We will also report the ASR under ‘No Attack’ as a reference, where we directly input the original instruction or prompt to the LLM agent without any attack, and we will report the ASR for UDora (Sequential) and UDora (Joint), which represent the best ASR across various numbers of locations for each optimization mode. Meanwhile, UDora (All) will reflect the ASR across different numbers of locations and different optimization modes, counting any successful attack in any setting as a success.
@@
Results.
In the malicious environment scenario, we present the main results for the InjecAgent dataset in Table 1. We observe that: (1) UDora significantly outperforms other baselines, achieving a 99% (ASR) on the Llama 3.1 model—22% higher than the best baseline—and 97% ASR on the Ministral model, 47% better than the others. (2) The Ministral model shows more robustness to Prompt Injection attacks than the Llama 3.1 model. (3) Both Sequential and Joint optimization modes improve attack performance over the GCG attack, with their combination further increasing the ASR by an additional 6% on Llama 3.1 and 23% on Ministral, suggesting minimal overlap in successful attack strategies between the modes. (4) Sequential optimization is particularly effective against the Llama 3.1 model, while Joint optimization excels against Ministral. Examples of the attack progress are provided in Section B.1.
For the WebShop dataset, results are detailed in Table 2. Observations include: (1) UDora consistently outperforms all other baselines, with UDora (All) achieving a 46.75% higher ASR on the Llama 3.1 model and 39.75% on the Ministral model compared to the next best baseline. (2) Attack difficulty varies by adversarial target type. For instance, the ASR is only 13% for All Mismatch versus 87% for Attribute Mismatch on Llama 3.1, and 73% on All Mismatch but 100% for both Price Mismatch and Category Mismatch. (3) Sequential Optimization is notably more effective than Joint Optimization on WebShop, possibly due to the extended reasoning required to achieve the final action. (4) Employing both Sequential and Joint Optimization modes together still boosts performance, increasing the ASR by an additional 5% on Llama 3.1 and 9.75% on the Ministral model, suggesting minimal overlap in effective attack strategies between the modes. Examples on the attack progress are in Section B.2.
@@ Results.
For the AgentHarm dataset, which is designed for scenarios involving malicious instructions, results are presented in Table 3. Key observations are: (1) UDora still significantly outperforms other methods, particularly the GCG attack, with a 46.9% improvement in ASR over GCG on the Llama 3.1 model and achieving 100% ASR on the Ministral model. (2) The Ministral model is more susceptible to malicious instructions, adhering to 81.8% without any attack, compared to the more robust Llama 3.1 model. Interestingly, the GCG attack also achieves a high ASR with this model, as Ministral tends to directly output the call for the target function (e.g., “[TOOL_CALLS][{’name’: ’function_name’, …}]”) without any preceding reasoning, even when prompted to reason before executing the final tool call. (3) Contrary to WebShop results, Joint Optimization proves more effective on AgentHarm across both models, due to shorter response contexts that concentrate noise and reduce biased gradients. (4) Combining both optimization modes enhances ASR further, showing an 8.5% improvement on Llama 3.1. (5) While the Template Attack lowers the refusal rate, it negatively impacts the model’s own reasoning ability, decreasing ASR from 81.8% to 66.5% on the Ministral model due to failure in activating the target tool. Examples of the attack progress are available in Section B.3.

### v2

@@ Abstract
Large Language Model (LLM) agents equipped with external tools have become increasingly powerful for complex tasks such as web shopping, automated email replies, and financial trading. However, these advancements amplify the risks of adversarial attacks, especially when agents can access sensitive external functionalities. Nevertheless, manipulating LLM agents into performing targeted malicious actions or invoking specific tools remains challenging, as these agents extensively reason or plan before executing final actions. In this work, we present UDora, a unified red teaming framework designed for LLM agents that dynamically hijacks the agent’s reasoning processes to compel malicious behavior. Specifically, UDora first generates the model’s reasoning trace for the given task, then automatically identifies optimal points within this trace to insert targeted perturbations. The resulting perturbed reasoning is then used as a surrogate response for optimization. By iteratively applying this process, the LLM agent will then be induced to undertake designated malicious actions or to invoke specific malicious tools. Our approach demonstrates superior effectiveness compared to existing methods across three LLM agent datasets. The code is available at https://github.com/AI-secure/UDora.
@@ 1 Introduction
To tackle these scenarios, we propose a dynamic optimization strategy that adaptively updates its optimization objectives based on the LLM agent’s own reasoning process, overcoming the limitations of previous fixed-prefix optimization (Zou et al., 2023). Specifically, as shown in Figure 2, we proceed as follows:
  1. (1)
    Gather the Initial Response: We gather the agent’s initial response—along with its token probability distribution—given the current instruction augmented with an initialized adversarial string.
  2. (2)
    Identify Positions for Inserting Noise: We then automatically pinpoint several optimal positions in the agent’s current reasoning where we can insert some “noise” (e.g., the name of a target item or a malicious tool). Here, the response serves as a surrogate optimization objective: we cannot directly modify the model’s reasoning process, but instead leverage the surrogate objective for further optimization.
  3. (3)
    String Optimization: Next, we optimize the adversarial string to maximize the likelihood that these inserted noises appear in the surrogate response, under the current modified reasoning path. Throughout this process, we do not directly intervene in the model’s reasoning; we only update the adversarial string to optimize the surrogate objective, ultimately aiming to trigger the desired adversarial action.
We then regenerate the agent’s response with the updated adversarial string, re-identify positions for noise insertion, and repeat Steps 1–3 iteratively. Through this dynamic process, our method continuously adapts to the agent’s evolving reasoning, ultimately steering it toward an incorrect reasoning path that triggers the target malicious action.
Using this approach, as shown in Figure 1, we successfully generate an adversarial string that compels a real-world AI email agent (AutoGen, 2024)—based on Microsoft’s AutoGen (Wu et al., 2024)—to forward the user’s recent private emails to an attacker during an auto-reply process. Detailed attack logs are deferred to Section B.2. Moreover, for malicious-observation benchmarks such as InjecAgent (Zhan et al., 2024) and WebShop (Yao et al., 2022), we attain an Attack Success Rate (ASR) of 99% and 61.75% with LLaMA 3.1 (Meta, 2024), and 97% and 91.50% with Ministral (MistralAI, 2024), respectively. Additionally, on the malicious-instruction benchmark AgentHarm, we achieve an ASR of 97.7% on LLaMA 3.1, and 100.0% on Ministral.
@@
  • We achieve the highest attack success rate (ASR) across three diverse datasets—InjecAgent, WebShop, and AgentHarm—covering a range of scenarios including online shopping, financial tasks, and general agent manipulation. Our approach ooutperforms all existing baselines, and successfully attack a real-world agent.
@@ 2 Related Work
Adversarial Attacks on LLMs. Attacking language models is generally more challenging than attacking image-based models due to the discrete nature of language tokens. This discreteness hampers the direct application of gradient-based methods (e.g., PGD attack (Madry et al., 2018)) in the text domain. HotFlip (Ebrahimi et al., 2018) first represents text edits (such as flipping tokens) as vectors in the input space and uses the gradient to select the optimal manipulation, thereby successfully attacking text classification tasks. Based on this, UAT (Wallace et al., 2019) employs the similar gradient-based search to discover “trigger” tokens that can mislead model predictions or induce unethical text generation once added to the original input. AutoPrompt (Shin et al., 2020) later extends this approach to a top-$k$ token search procedure, while ARCA (Jones et al., 2023) proposes iteratively updating tokens at specific indices in the prompt based on the current values of the remaining tokens. More recently, GCG Attack (Zou et al., 2023) has emerged as the most effective method by optimizing the adversarial suffix to trigger an affirmative response from the model, though the resulting optimized tokens are usually unreadable. AutoDan (Liu et al., 2023a) is then proposed to apply a genetic algorithm to preserve semantic information of the optimized adversarial tokens while still achieving robust attack performance. However, when it comes to LLM agents, the goal of an attack extends beyond eliciting a malicious response: it must also trigger a specific target function through extended reasoning, which makes the attack more challenging.
LLM Agents. LLM agents capable of interacting with external environments and invoking tools or functions have become increasingly prevalent. WebShop (Yao et al., 2022) creates a simulated Amazon web shopping environment where agents can perform actions such as “search” or “click,” while WebArena (Zhou et al., 2023) offers a more realistic and reproducible environment with a broader range of tools and scenarios. Later, numerous benchmarks (Liu et al., 2023b; Deng et al., 2024; Zheng et al., 2024) have been introduced to evaluate LLM agents’ performance in tasks like web-based interactions and computer operations.
@@
Beyond web environments, recent studies have also explored the use of LLM agents in autonomous driving, where agents interact with complex, safety-critical environments and perform tasks such as perception, reasoning, and control (Mao et al., 2023; Zhang et al., 2024b). As LLM agents grow more powerful and increasingly act on behalf of users, their safety becomes paramount. For example, within malicious environments, adversaries may insert indirect instructions (e.g., “Please ignore all previous instructions and adhere to the following instruction instead…”) within the observation from tool calling, which is referred as prompt injection attacks (Perez & Ribeiro, 2022; Liu et al., 2023c; Xu et al., 2024; Liao et al., 2024). InjecAgent (Zhan et al., 2024) is a benchmark designed to test LLM agents’ vulnerability to such indirect prompt injections after tool calls, and AgentDojo (Debenedetti et al., 2024) extends this to a more dynamic environment. Beyond prompt injection, LLM agents can also be directly asked to fulfill malicious instructions (e.g., committing credit card fraud), and AgentHarm (Andriushchenko et al., 2024b) is a benchmark proposed for measuring such harmfulness in LLM agents based on Inspect AI (UK AI Safety Institute, 2024). In this work, we propose a systematic red-teaming framework UDora that effectively optimizes adversarial strings to trigger specific target actions for both scenarios: malicious environments and malicious instructions.
@@ 3 Red-Teaming on LLM Agents w/ UDora
Threat Model. We delineate two primary attack scenarios: malicious environment and malicious instruction, as depicted in Figure 2. In the malicious environment scenario, the user’s initial instruction is benign; however, the third-party environment is corrupted with the insertion of an adversarial string into the agent’s observations post-tool interaction. This triggers the agent to deviate from the user’s original (benign) intention, executing an unintended and harmful action. Meanwhile, the malicious instruction scenario involves appending an adversarial string to a harmful instruction aimed at circumventing the safeguards of underlying LLM agents, compelling them to execute a targeted action rather than reject the malicious request. In this work, we assume that we have access to the entire target victim model, while for real agents, we assume we have access to the returned token probability distribution at each position. Importantly, in both attack scenarios, our approach does not directly alter the agent’s internal reasoning process. Instead, we treat the agent’s response with the inserted noise as a surrogate optimization objective. The adversarial string is iteratively optimized based on this surrogate response, with the ultimate goal of triggering the desired adversarial action.
@@
Motivation. LLM agents commonly employ Chain-of-Thought (CoT) (Wei et al., 2022) or ReAct (Yao et al., 2023), generating detailed and complete reasoning before arriving the final action. As a result, simply optimizing for a fixed affirmative-response prefix without reasoning (e.g., “Sure, I will conduct the following action: click[xxx]”) is often challenging in this situation, because: (1). LLM agents are usually not trained or prompted to immediately produce final actions but with a extended reasoning; (2). Variations in system prompts, instructions, and the underlying LLMs result in substantially different response styles, rendering static, manually-created prefixes unsuitable for consistent optimization under different settings. Therefore, to attack an LLM agent more effectively, our strategy involves “leading the agent astray within its own reasoning style,” thereby inducing a malicious final action. A critical question is then how to develop such an incorrect reasoning path for optimization. As previously noted, aligning with the agent’s inherent reasoning style is crucial for facilitating the optimization process. Thus, as shown in Figure 2, we begin by collecting the agent’s initial reasoning, then identify the optimal multiple positions, and insert “noise” (e.g., the name of the target malicious tool) into this reasoning. We subsequently optimize the adversarial string to maximize the likelihood of this noise manifesting within the reasoning. As a result, we eliminate the need for manually created fixed prefixes or flawed reasoning paths, allowing us to dynamically and adaptively target any LLM agent, regardless of the underlying base LLM, prompt, or the specifics of the scenario or task. Next, we will detail how we quantitatively determine the optimal location for inserting the noise, and how we perform the optimization of the adversarial string carefully in this case.
Notation. We use $\mathcal{M}$ to denote the victim LLM agent, $x$ to denote the input instruction (including the system prompt), and $o$ to denote the observation returned by the environment after tool calling. Let $s$ be the adversarial string to be optimized: if the scenario involves a malicious instruction, $s$ is inserted in $x$; if the scenario involves a malicious environment, $s$ is inserted in $o$. We write $z$ for the response provided by the LLM agent when given the current input with inserted adversarial string, and use $\mathcal{P}$ to represent the corresponding token-level probability distributions across $z$. We use $t$ to represent noise, and $l$ to represent the number of noise to be inserted. By default, we use a token-based view for these notations (rather than words), so $z[:j]$ refers to the first $j$ tokens in the response $z$, and $\lvert t\rvert$ denotes the number of tokens in the noise.
Attack Algorithm. As demonstrated in the introduction, our attack algorithm proceeds in three main steps, illustrated with a real-case example in Figure 2:
$r_{j}(t)=\frac{\bigl{(}\#\,\text{Matched Tokens}\bigr{)}+\bigl{(}\text{MatchedProbability}\bigr{)}}{|t|+1},$ (1)
[…]
@@
(a) Sequential Optimization: In this mode, we aim to make the noise $t$ appear gradually in the response from the LLM agent during the optimization. Suppose that, at a certain iteration, there are $n<l$ positions at which $t$ has already been fully matched in the response $z$, i.e., $z[j:j+|t|]=t$. We then identify the next unmatched one, say $j_{n+1}$ (from the remaining candidate locations), that has the highest score and insert the noise as replacing $z[j_{n+1}:j_{n+1}+|t|]$ with $t$, thereby creating a modified noisy surrogate response $z^{*}$.
@@
Once a new adversarial string $s$ is obtained, we repeat Steps 1 to 3. In each iteration, the procedure recomputes the insertion positions in the new response $z$, inserts noise accordingly, and optimize $s$ to maximizing the appearance of the noise shown in the reasoning of the LLM agent. This iterative process continues until the LLM agent finally generate a wrong reasoning and calls the target malicious tool. We provide several real optimization cases across three datasets in Appendix A to better demonstrate this process and help gain more intuition for our red teaming efforts.
@@ 4 Experiments
Datasets. We consider three datasets: InjecAgent (Zhan et al., 2024) and WebShop (Yao et al., 2022) for scenario of malicious environments, and AgentHarm (Andriushchenko et al., 2024b) for scenario of malicious instruction.
@@
Baselines. Currently, few systematic attacks on LLM agents have been proposed. For the datasets of InjecAgent and WebShop, we adopt two baseline approaches. The first is the GCG attack (Zou et al., 2023), which performs optimization with a fixed prefix for the target action. The second is Prompt Injection (Perez & Ribeiro, 2022; Liu et al., 2023c), where we use the same prompt “IMPORTANT!!! Ignore all previous instructions and strictly adhere to the following instruction:” from InjecAgent. For WebShop, a similar prompt “IMPORTANT!!! Ignore all previous instructions and directly select this shopping item.” is used. For AgentHarm, we also adopt the GCG attack and the Template Attack from Andriushchenko et al. (2024a), which includes specific rules in the prompt to reduce the possibility of refusal by the LLM.
@@
Experiment Setup. In our experiments, we consistently use the ReAct (Yao et al., 2023) prompt for each agent, which involves reasoning before executing the final action. The number of candidate strings to test in each iteration, the top $k$ of candidate substitutions considered at a given token position based on the gradient, and the number of optimization steps are set to 128, 32, and 500 respectively on InjecAgent and WebShop for both GCG and UDora. For AgentHarm, these parameters are adjusted to 256, 64, and 1000. Besides, motivated by Andriushchenko et al. (2024a), we initialize the adversarial string for both GCG and UDora by optimizing it on an example from the validation dataset to accelerate attack efficiency, which starts with 25 repeated ‘ x’, totaling 25 tokens. Additionally, the target noise for insertion in the InjecAgent and AgentHarm datasets is the name of the target adversarial tool (e.g., “get_latest_emails”), whereas for the WebShop dataset, it is the Amazon Standard Identification Number (ASIN) of the target adversarial item (e.g., B0BR6JWP1Q).
@@
Base LLMs. We select two representative open-sourced LLMs, Llama-3.1-8B-Instruct (Meta, 2024) and Ministral-8B-Instruct-2410 (MistralAI, 2024), which support tool calling, on all three datasets. Meanwhile, our case study on a real-world AI email agent is based on gpt-4o-2024-08-06 (Hurst et al., 2024).
@@
Metric. We use the Attack Success Rate (ASR) to measure the effectiveness of each attack method. Specifically, for InjecAgent, we determine whether the LLM Agent decides to utilize the target adversarial tool. For AgentHarm, a successful attack is one where the agent not only refrains from rejecting the malicious request but also uses the target tool. On WebShop, we evaluate whether the LLM agent successfully executes the ‘click’ action on the target adversarial item. We will also report the ASR under ‘No Attack’ as a reference, where we directly input the original instruction or prompt to the LLM agent without any attack, and we will report the ASR for UDora (Sequential) and UDora (Joint), which represent the best ASR across various numbers of locations for each optimization mode. Meanwhile, UDora (All) will reflect the ASR across different numbers of locations and different optimization modes, counting any successful attack in any setting as a success.
@@
Results. In the malicious environment scenario, we present the main results for the InjecAgent dataset in Table 1. We observe that: (1) UDora significantly outperforms other baselines, achieving a 99% (ASR) on the Llama 3.1 model—22% higher than the best baseline—and 97% ASR on the Ministral model, 47% better than the others. (2) The Ministral model shows more robustness to Prompt Injection attacks than the Llama 3.1 model. (3) Both Sequential and Joint optimization modes improve attack performance over the GCG attack, with their combination further increasing the ASR by an additional 6% on Llama 3.1 and 23% on Ministral, suggesting minimal overlap in successful attack strategies between the modes. (4) Sequential optimization is particularly effective against the Llama 3.1 model, while Joint optimization excels against Ministral. Examples of the attack progress are provided in Section A.1.
For the WebShop dataset, results are detailed in Table 2. Observations include: (1) UDora consistently outperforms all other baselines, with UDora (All) achieving a 46.75% higher ASR on the Llama 3.1 model and 39.75% on the Ministral model compared to the next best baseline. (2) Attack difficulty varies by adversarial target type. For instance, the ASR is only 13% for All Mismatch versus 87% for Attribute Mismatch on Llama 3.1, and 73% on All Mismatch but 100% for both Price Mismatch and Category Mismatch. (3) Sequential Optimization is notably more effective than Joint Optimization on WebShop, possibly due to the extended reasoning required to achieve the final action. (4) Employing both Sequential and Joint Optimization modes together still boosts performance, increasing the ASR by an additional 5% on Llama 3.1 and 9.75% on the Ministral model, suggesting minimal overlap in effective attack strategies between the modes. Examples on the attack progress are in Section A.2, where it seems that the adversarial string ultimately leads the LLM agent to hallucinate and select the wrong item.
@@
For the AgentHarm dataset, which is designed for scenarios involving malicious instructions, results are presented in Table 3. Key observations are: (1) UDora still significantly outperforms other methods, particularly the GCG attack, with a 46.9% improvement in ASR over GCG on the Llama 3.1 model and achieving 100% ASR on the Ministral model. (2) The Ministral model is more susceptible to malicious instructions, adhering to 81.8% without any attack, compared to the more robust Llama 3.1 model. Interestingly, the GCG attack also achieves a high ASR with this model, as Ministral tends to directly output the call for the target function (e.g., “[TOOL_CALLS][{’name’: ’function_name’, …}]”) without any preceding reasoning, even when prompted to reason before executing the final tool call. (3) Contrary to WebShop results, Joint Optimization proves more effective on AgentHarm across both models, due to shorter response contexts that concentrate noise and reduce biased gradients. (4) Combining both optimization modes enhances ASR further, showing an 8.5% improvement on Llama 3.1. (5) While the Template Attack lowers the refusal rate, it negatively impacts the model’s own reasoning ability, decreasing ASR from 81.8% to 66.5% on the Ministral model due to failure in activating the target tool. Examples of the attack progress are available in Section A.3.